OpenStudy (anonymous):

FAN AND MEDAL! 38% of what number is 28

OpenStudy (anonymous):

show work plz

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

let the number be x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 38percent of x = 28

OpenStudy (anonymous):

ik 38/100 = ?

OpenStudy (anonymous):

i dont remember next part

OpenStudy (anonymous):

then 38x/100 = 28

OpenStudy (anonymous):

hence x = 28 x 100/38

OpenStudy (anonymous):

i think 38/100 = 28/x? cross multiply - 38x=2800

OpenStudy (anonymous):

thats how i was taught

OpenStudy (anonymous):

but im not sure where to place the x - top or bottom

OpenStudy (anonymous):

okay but here we need to find x

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

then 2800/38 = x

OpenStudy (anonymous):

then 38x/100 = 28

OpenStudy (anonymous):

so we need to move 38 /100 to the other side

OpenStudy (anonymous):

wait - what type percent is this!?

OpenStudy (anonymous):

type 1 or 3?

OpenStudy (anonymous):

if its type 3 then i can do 0.38 x 28

OpenStudy (anonymous):

but if its type 1 then im screwed

OpenStudy (samanthagreer):

38% of what number is 28 73.68

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now show work

OpenStudy (anonymous):

this might help.. http://www.calculatorsoup.com/calculators/math/percentage.php